OpenStudy (anonymous):

The main tank has a radius of 70 feet. What is the volume of the quarter-sphered sized tank? Round your answer to the nearest whole number. You must explain your answer using words, and you must show all work and calculations to receive credit.

OpenStudy (popitree):

volume of sphere=> 4/3\[4/3\pi r ^{3}\]; hence 1 quarter = \[1/3\pi r ^{3}\] hence the ans is 359006.67

OpenStudy (popitree):

ignore 1st 4/3..typing mistake

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

how do I do this problem??